# Renormalization Group Evolution for In-medium Energy Correlators

### v1

@@ 6.2 Results and Discussion for the Energy Correlator in $p$-Pb and O-O
We explore the theoretical uncertainty on the example of $p_{T}=100$ GeV jets. The band of Figure 9 corresponds to a variation of the the LPM scale by a factor of 2: $\sqrt{2p_{\rm jet}/L_{\rm eff}}<\mu_{LPM}<2\sqrt{2p_{\rm jet}/L_{\rm eff}}$. The uncertainty band from the variation of this semi-hard scale is reasonably under control. Besides, there are many soft parameters used in above calculation related to the medium properties, including
  • The initial time $\tau_{0}$ when the jet parton starts to couple to the medium. The uncertainty comes from the lack of knowledge of when the color charges of the QGP are fully produced from the initial scattering. s It is varied from $0.1$ fm/$c$ to $0.5$ fm/$c$.
@@ Appendix B On the soft radiation outside of the jet cone
The opacity-one correction to soft radiation outside of the jet cone is given by
$\mathcal{H}_{ab\rightarrow qX}\otimes\mathcal{J}_{q}^{(1)}\approx$ $\int\frac{dx}{x}\mathcal{H}_{ab\rightarrow qX}\left(\frac{p_{T}}{x}\right)\left[\frac{dN^{N=1}_{q\rightarrow q+g}}{dxd^{2}\mathbf{k}}\right]_{+}\Theta_{{\rm anti}-k_{T}}^{>R}d^{2}\mathbf{k}\;.$ (140)
First, we use the soft gluon approximation for the splitting function and consider only the transverse integral. Next, we shift $\mathbf{q}\rightarrow\mathbf{q}+\mathbf{k}$, define $t=|\mathbf{q}|^{2}/|\mathbf{k}|^{2}$ and reorganize the integral into
@@
which relates the hard parton spectrum between different scales. Choosing the other boundary as the medium screening scale $\mu^{2}=m_{\rm eff}^{2}$, where no medium-induced radiations is included, we have an analytic solution under the soft-gluon approximation

### v2

@@ 6.2 Results and Discussion for the Energy Correlator in $p$-Pb and O-O
We explore the theoretical uncertainty on the example of $p_{T}=100$ GeV jets. The band of Figure 9 corresponds to a variation of the the LPM scale up and down by a factor of 2. The uncertainty band from the variation of this semi-hard scale is reasonably under control. Besides, there are many soft parameters used in above calculation related to the medium properties, including
  • The initial time $\tau_{0}$ when the jet parton starts to couple to the medium. The uncertainty comes from the lack of knowledge of when the color charges of the QGP are fully produced from the initial scattering. It is varied from $0.1$ fm/$c$ to $0.5$ fm/$c$.
@@ Appendix B On the soft radiation outside of the jet cone
Following Eq. 5.2, the opacity-one correction to soft radiation outside of the jet cone is given by
$\mathcal{H}_{ab\rightarrow qX}\otimes\mathcal{J}_{q}^{(1)>}\approx$ $\int\frac{dx}{x}\mathcal{H}_{ab\rightarrow qX}\left(\frac{p_{T}}{x}\right)\int d^{2}\mathbf{k}\frac{dN^{N=1}_{q\rightarrow q+g}}{dxd^{2}\mathbf{k}}\Theta_{{\rm anti}-k_{T}}^{>R}\;,$ (139)
$=$ $\int\frac{dx}{x}\mathcal{H}_{ab\rightarrow qX}\left(\frac{p_{T}}{x}\right)I_{qq,R}(x)\;.$ (140)
First, we use the soft gluon approximation for the splitting function and complete the transverse integral to arrive at the kernel $I_{qq,R}(x)$. In computing $I_{qq,R}(x)$, we shift $\mathbf{q}\rightarrow\mathbf{q}+\mathbf{k}$, define $t=|\mathbf{q}|^{2}/|\mathbf{k}|^{2}$ and reorganize the integral into
@@
which relates the hard parton spectrum between different scales. Choosing the other boundary as the medium screening scale $\mu^{2}=m_{\rm eff}^{2}$, where no medium-induced radiation is included, we have an analytic solution under the soft-gluon approximation